OpenStudy (anonymous):

Let A be the area of a circle with radius r. If , dr/dt=2, find dA/dt when r=1

OpenStudy (anonymous):

A = pi*r^2 dA/dt = 2 * pi * r * dr/dt dr/dt = 2, r = 1: dA/dt = 2 * pi * 2 = 4 * pi